Let a = 25 degrees and b = 155 degrees. Which is a + b, in radians?

A. 4π/3
B. π/3
C. 0.8π
D. π
1
Expert's answer
2013-09-06T13:07:47-0400
a+b=25+155=180 (degrees) or π (in radians).
Answer: D. π.